Q8. A manufacture is considering a switch from manufacturers’ representatives to an internal sales force. Manufacturers’ reps are paid 8.5% commission and incur $625,000 in fixed costs; while an internal sale forced has fixed costs projected at $1,950,000 and would receive a 3% commission.

At what sales volume would the manufacture be indifferent between the two alternatives?  
Answer #1
Lest the sales volume be $x
625000+0.085x = 1950000+0.03x
0.085x-0.03x = 1950000-625000
0.055x = 1325000
x = 1325000/0.055= $24090909
Required sales volume = $24090909